### v2

# Structure-Aware Compound-Protein Affinity Prediction via Graph Neural Network with Group Lasso Regularization thanks: Code available: https://github.com/FrancisShizy/SAGGLR

Abstract
Artificial intelligence accelerates drug discovery via learning drug structure and molecular representation. Explainable artificial intelligence (XAI) approaches are increasingly applied in such areas to rationalize drug property prediction. However, building end-to-end explainable models for structure-activity relationship (SAR) modeling for compound property prediction faces many challenges, such as the limited number of compound-protein interaction activity data for specific protein targets, and plenty of subtle changes in molecular configuration sites significantly affecting molecular properties. Thus, optimally leveraging structural and property information and identifying key moieties related to compound–protein affinity for specific targets is essential. In this context, we exploit pairs of molecules with activity cliffs that share scaffolds but differ at substituent sites, characterized by large potency differences for specific protein targets. We propose a framework by implementing graph neural networks (GNNs) to leverage property and structure information from activity cliff pairs to predict compound-protein affinity (i.e., half maximal inhibitory concentration, IC50). To enhance model performance and explainability, we train GNNs with structure-aware loss functions using group lasso and sparse group lasso regularizations, which prune and highlight molecular subgraphs relevant to activity differences. We applied this framework to activity cliff data of molecules targeting three proto-oncogene tyrosine-protein kinase Src proteins (PDB IDs: 1O42, 2H8H, 4MXO). Our approach improved property prediction by integrating common and uncommon node information with sparse group lasso, as reflected in reduced root mean squared error (RMSE) and improved Pearson’s correlation coefficient (PCC). Applying regularizations also enhances feature attribution for GNN by boosting graph-level global direction scores and improving atom-level coloring accuracy. These advances strengthen model interpretability in drug discovery pipelines, particularly for identifying critical molecular substructures in lead optimization.
Keywords Drug Discovery, Drug Property Prediction, Graph Neural Network, Regularization
1 Introduction
Artificial intelligence accelerates drug discovery with methods such as structure-based virtual screening and quantitative structure-activity relationship (QSAR) modeling to learn drug structure information and molecular representation [1]. These models are typically defined as regression or classification models that establish relationships between molecular structure and properties for downstream prediction [2]. Deep learning models such as graph neural networks (GNN) [3] and message-passing neural network [4] (MPNN) models learn and update atom- and edge-level information of molecular graphs. They can enable end-to-end tasks such as predicting drug properties by entering the chemical structure of the molecules using the simplified molecular input line entry system (SMILES) [5]. In addition, explainable artificial intelligence (XAI) technologies have demonstrated their ability to accelerate drug discovery and the potential in designing small molecule and protein-based drugs [6]. Specifically, they can be used to predict drug-protein interactions for cancer [7], determine drug targets in neurological disorders [8], identify potential leads in in silico screening for COVID-19 [9], and predict the potency of compound for cardiovascular disease [10].
Despite these promising potentials in drug discovery, applying deep learning models to predict structure-activity relationships (SAR) faces several challenges. One challenge is data imbalance, which often occurs as chemical databases such as ChEMBL [11] contain a much larger number of inactive compounds than active compounds for a given protein target. Furthermore, numerous subtle alterations in molecular configuration sites can significantly impact molecular properties, increasing model complexity and affecting model capacity for end-to-end tasks such as drug property prediction. XAI methods have been increasingly applied in such areas to interpret drug property prediction results. For example, Rao et al. (2022) presented experimental results demonstrating that XAI methods can provide reliable and informative explanations to chemists in identifying key substructures [12]. However, many previous deep learning approaches that claim their capacity to explain GNN models for graph classification [13] [14] [15] often highlight inconsistent molecular substructures, leading to unstable model explainability for drug property prediction.
In this paper, we developed a novel framework by leveraging the matched molecular pair(MMP)-based cliff in the training to discern key moiety structures related to the compound-protein affinity. We also modified the traditional loss function by adding regularization terms using group lasso and sparse group lasso to improve explainability. We mainly applied our framework on predicting compound affinity for Src [16], an important protein kinase that is important for many diseases such as Alzheimer’s disease [17] and cancers [18]. Our study leads to the findings below: (i) Deep learning model performance for drug property prediction is improved by using graph neural networks by capturing and integrating common and uncommon substructure information of activity-cliffs pairs in the model. The improvement also takes advantage of the loss functions with regularization to prune and highlight molecular subgraphs. (ii) Applying regularization methods such as group lasso and sparse group lasso enhance feature attribution by boosting global direction scores and atom-level accuracy in atom coloring prediction. They provide stable subgraph-level model explainability for molecular property prediction. (iii) Our approach can enhance the interpretability of models in drug discovery and virtual screening pipelines, particularly in investigating important molecular substructures in lead optimization. We introduce the rationales of our framework below in details.
Drug-protein interaction (DPI) modeling tasks, including compound-protein affinity prediction and functional group-based drug property prediction and classification, are important research areas for traditional SAR models and graph-level tasks of deep learning models, such as GNN models. These models have been applied to preclinical drug discovery, virtual screening, and lead optimization [2] [19]. Instead of directly inputting imbalanced drug-protein interaction data into such models, one strategy is to combine molecular activity information such as half-maximal inhibitory concentration (IC50) with structure information (e.g., scaffolds and motifs) to integrate datasets. The loss functions for the models are defined based on molecular similarity and potency difference criteria. Activity cliffs (ACs) are generally defined as pairs or groups of structurally similar compounds that are active against the same target but have large differences in potency [20]. A matched molecular pair (MMP) is a pair of compounds that share a common core structure and are distinguished by substituents at a single site [21]. MMP-based AC (MMP-cliff) can be exploited to capture a large difference in potency between the participating compounds by applying the maximum common substructure (MCS) formalism [22] to calculate the molecular scaffold between pairs of compounds binding to a specific target (Figure 1). For two compounds to be considered as sharing a molecular scaffold, the shared substructure must constitute at least a specified fraction of each molecule’s framework (i.e., percentage thresholds of shared atoms between the matched pairs). Thus, such activity-cliff data can provide critical information regarding key structural components that determine the drastic difference in MMPs. They can then be applied to predict drug-protein binding affinities and provide topology-specific model explainability for drug property differences for specific protein targets. In many previous studies [21][22][23], models were built by training on compound-protein interaction datasets, including various targeting proteins downloaded from cheminformatics databases, to predict compound-protein affinity. While these models achieved reasonable average accuracy over a large number of targets, the accuracy on individual protein targets is often unsatisfactory. By exploiting the information offered by MMP-cliff, we can focus our training on data for a specific protein target.
Figure 1: An illustration for an example pair of molecules targeting Src kinase 1O42 with activity cliffs ($\Delta pK_{i}=3.85$). The paired two molecules share common and uncommon substructures. Common substructures consist of uncolored nodes (atoms) and edges (bonds), and uncommon substituents consist of colored nodes and edges (blue & red circles for sites in the pair)
Besides tasks like drug property prediction, many previous XAI models have explored node- and edge-level tasks in order to identify crucial molecular substructures [24] [12] for DPI. As an analytical interpretation of SAR models, color coding aims to assign an importance value for every input feature, such as an atom or a bond in a molecular graph. Such importance values are often visualized through atom or bond coloring, where the structural patterns that drive a prediction are highlighted on the molecular representation of the compound of interest. Feature attribution methods can identify and highlight vital substructures that determine predicted property change by estimating the contribution of each atom in the molecular graphs [25]. The consistency of the highlighted moieties, combined with expert background knowledge, is expected to enhance the understanding of machine learning models in drug design  [22]. After exploiting various feature attribution techniques (e.g., Gradient $\times$ Input (DeepLIFT)[26], Gradient-weighted Class Activation Mapping (Grad-CAM) [27]) to extract the node information for the structural difference of compound pairs, we can assess feature attribution performance to estimate and compare the accuracy of model explainability under different model training settings designed for extracting various node and edge information in molecular graphs.
From the perspective of graph learning, the nodes and edges correspond to scaffolds and decorations, which can be regarded as distinct subgraphs within an overall molecular graph. We aim to train an explainable deep learning model that can identify useful features for drug property prediction, which are learned by keeping significant information from vital nodes in a natural group structure (i.e., subgraphs) and pruning those not significant. If we view such features as “covariates” for drug property prediction, the various forms of group lasso penalty are designed for such situations that have all coefficients within a group become nonzero (or zero) simultaneously. Activity cliffs are generally defined as pairs or groups of structurally similar compounds that are active against the same target but have large differences in potency. Therefore, based on the concept of active-cliff pairs, it is intuitive to categorize nodes in molecular graphs into common (i.e., scaffolds) and uncommon (i.e., decorations) groups. Group lasso can then be used to prune and select node-level information by either selecting or discarding entire groups of features together to help explain the activity cliff.
Overall, our framework aims to maximally utilize the information contained in the MMP-cliffs in improving model accuracy as well as model explainability for drug property prediction and discovery. Unlike many current works that develop models for the entire drug-protein database, we focus on predicting drugs targeting individual protein families.
2 Methods
In this project, we propose the GNN/MPNN framework as shown in Figure 2. Specifically, we first conducted drug property database screening and active pairs generation. Next, we trained deep learning models such as GNN for compound-protein affinity prediction to represent molecular scaffolds with common and uncommon nodes and extract molecular structure information. We also identified important molecular substructures for such molecular properties using MPNN and GNN models with and without regularization techniques like group lasso and sparse group lasso [28]. Furthermore, we evaluated and compared the performance of different models with various loss function settings. We then assessed feature attribution performance for model explainability using graph-level metrics and atom-level accuracy in atom coloring prediction.
Figure 2: Model structure illustration. Considering a pair of compounds $c_{i}$ and $c_{j}$ that share a scaffold in the red circle and decorations in the blue circle, GNN and MPNN were applied to learn latent node representations for both common and uncommon nodes. Such node-level information was aggregated to predict graph-level drug-protein binding affinity, and then the mean squared errors for predicted and experimental ones were calculated. Additionally, masking functions for common and uncommon nodes and readout functions were applied to normalize the information, and multilayer perceptrons (MLPs) were combined with group lasso and sparse group lasso to create subgraph node loss for both common and uncommon nodes. The two loss functions were minimized during the training process. "CN" for common nodes and structure, "UCN" for uncommon nodes and structure, "N" for both common and uncommon nodes. Regularization methods were also applied to prune and select node-level information in activity cliff pairs. Group lasso considers only group-level sparsity (nodes within the same subgroup having the same color depth representing the same weights after pruning), and sparse group lasso considers both group-level and within-group sparsity (nodes within the same subgroup having different color depth representing various weights after pruning).
2.1 Data preparation
The activity-cliff pair data were generated based on the MCS of closely related compounds featuring property cliffs for model training. Specifically, activity cliff pairs were defined as compound pairs within one or more congeneric series that share a molecular scaffold and exhibit at least 1 log unit activity difference of small molecule activity data (e.g., half maximal inhibitory concentration, IC50). In this study, the IC50 values for the target proteins and associated small molecules derive from protein-based assays. The IC50 data for the Src tyrosine kinase family proteins and associated small molecule activity for generating activity-cliff pairs are originally downloaded and constructed from the BindingDB validation datasets [29]. We created the matched molecular pairs (MMP) of compounds targeting each kinase in the Src family respectively by the FMCS algorithm in the RDKit rdFMCS module [22]. We then screened compounds based on some structural similarity criteria to expand the current datasets of the Src tyrosine kinase family. In addition, only protein targets with at least 50 compound pairs in the training set were kept. Three tyrosine-protein kinase Src proteins and related compounds were selected from the 25 tyrosine-protein kinase proteins. There are 1,377 activity-cliff pairs consisting of 90 compounds targeting the kinase with PDB ID 1O42, 480 pairs consisting of 63 compounds targeting 2H8H, and 10170 pairs consisting of 216 compounds targeting 4MXO [30]. We split the compound pair dataset into a 70% training set, 10 % validation set and a 20% testing set for each kinase.
2.2 Loss function and regularization
Model framework
Figure 2 illustrates the model training including molecular representation learning and loss function settings. For any pair $k$ of compounds $i$ and $j$, corresponding to molecular graphs $c_{i},c_{j}\in C$, where $C$ is the graph set of all compound pairs, and drug properties (i.e., IC50) $y_{i},y_{j}\in\mathbb{R}$. After generating such activity-cliff pairs, the node and edge information was fed into the GNN and following layers, including (i) two linear layers for node and edge embedding respectively, (ii) followed by three convolutional layers (i.e., NNConv layers with aggregation method), each NNConv layer applies a neural network defined by a Sequential module containing a single linear transformation, (iii) three batch normalization and three ReLU activation layers, (iv) two linear layers that transform the feature dimension of attributions for common and uncommon nodes respectively, (v) and a linear layer to combine such attributions and an outcome layer for affinity prediction. We used early stopping during model training to mitigate overfitting to avoid possible overestimation of the algorithm’s performance.
Loss function design
Previous research [22][31] assumed that common substructures are considered neutral while uncommon structures and their latent space determine the drug property difference. They emphasized the uncommon nodes causing the activity change and proposed a loss scheme focusing on uncommon structural motifs. However, the structural integrity and connections between skeletons and motifs within molecules may collectively influence the drug properties. We thus applied graph representation learning approaches to learn the latent space of both common and uncommon atom information from molecular substructures using the designed loss function settings. A node loss function was specifically designed to capture common skeletons and uncommon node features causing the activity change during model training. We also defined the main loss function as the mean squared error (MSE) to minimize the difference between observed and predicted binding affinities for compound pairs during training. Furthermore, we utilized different regularization methods, including group lasso and sparse group lasso penalties on these loss functions, to prune and highlight important molecular substructures. Group Lasso is used when features can be naturally grouped into predefined groups, and it selects or discards entire groups of features together rather than individual features [32]. Sparse group lasso extends group lasso by allowing for both group-level sparsity and within-group sparsity [28]. This means it can further add flexibility to prune and select node-level information for a subset of individual features given the common and uncommon subgraph groups within activity-cliff pairs.
Basic loss functions for model training are:
$L_{\text{MSE}}(m_{k})=\sum_{k\in(i,j)}\bigg\|y_{k}-\hat{y}_{k}\bigg\|^{2}$
$L_{\text{N}}(m_{k})=\bigg\|\sum_{k\in(i,j)}\text{MLP}\bigg(R_{k}\big(F_{k}(h_{k})\big)\bigg)-y_{k}\bigg\|^{2}$
where $m_{k}$ is the active-cliff pair of compounds $(i,j)$ corresponding to their drug affinities $(y_{i},y_{j})$ and predicted values $(\hat{y}_{i},\hat{y}_{j})$, $\text{MLP}(\cdot)$ is a multi-layer perception with linear activation, $R_{k}(\cdot)$ is a mean readout function over nodes, $h_{k}$ is the latent node representation of compounds $i$ and $j$ after GNN representation learning, and $F_{k}(\cdot)$ is a masking function over nodes for compound $i$ and $j$ in the context of pair $k$ including the uncommon node masking and the common node masking that determines the node loss $L_{N}$ consists of uncommon structure loss ($L_{UCN}$) and common structure loss ($L_{CN}$).
For our framework, we also include additional Lasso terms [28] in our loss functions as described below:
(i) Apply group lasso:
$L_{\text{MSE}}+L_{\text{N}}+\lambda\bigg[\sqrt{p_{\text{CN}}}\bigg\|\beta^{(\text{CN})}\bigg\|_{2}+\sqrt{p_{\text{UCN}}}\bigg\|\beta^{(\text{UCN})}\bigg\|_{2}\bigg]$
where $\beta^{\text{CN}}$ and $\beta^{\text{UCN}}$ are the parameters for the MLP layers for pruning the common and uncommon node information, and $p_{(\cdot)}$ is the number of covariates in subgraphs of common nodes ($p_{\text{CN}}$) and uncommon nodes ($p_{\text{UCN}}$).
(ii) Apply sparse group lasso:
$L_{\text{MSE}}+L_{\text{N}}+(1-\alpha)\lambda\bigg[\sqrt{p_{(\text{CN})}}\bigg\|\beta^{(\text{CN})}\bigg\|_{2}+\sqrt{p_{(\text{UCN})}}\bigg\|\beta^{(\text{UCN})}\bigg\|_{2}\bigg]+\alpha\lambda\bigg\|\beta\bigg\|_{1}$
where $p_{(\cdot)}$ is the number of covariates in each group.
2.3 Feature attribution
Attribution is an approach to interpretability which highlights the input feature dimensions influential to a neural network’s prediction [33]. An attribution map, $G_{A}=(v_{A},e_{A})$, is generated by an attribution method $A$ taking a model $M$ and a graph $G$. The attribution map consists of node and edge weights, $v_{A},e_{A}\in\mathbb{R}$ respectively, which are relevant for predicting graph property $y$. These weights can be visualized as a heatmap overlaid on the graph, providing insights into the contributions of different nodes and edges to the prediction.
Feature attribution approaches can help explain which parts of the provided inputs are considered relevant by the underlying supervised learning method for a specific prediction [22]. In the context of drug discovery and molecular design, feature attribution techniques for XAI models typically involve the coloring of the molecular graphs by producing a real number of coloring for each atom in the graph [25]. Coloring techniques aim to assign an importance value to each input feature, such as an atom or bond in a molecular graph. These importance values are typically displayed as gradient colorings on atoms or bonds, emphasizing the structural features that drive predictions for the compound’s molecular representation. Jimenea L. et al. introduce the MCS algorithm to determine proxy ground-truth atom-level color labels for the considered sets [22]. We used this method to determine ground-truth atom-level feature attribution labels via the concept of activity cliffs. Each identified pair constitutes an activity cliff, and we assumed that the observed potency difference for the two compounds in a pair results from the structural variations of motifs and their connection with the common scaffolds. Corresponding to the influence of that substructural feature on the predicted property, we labeled molecular subgraphs as negatively influenced and positively influenced nodes with different colors. Features in the more active compound were expected to receive positive attribution, whereas features in the less active compound received negative attribution. Consequently, atomic labels are assigned according to the activity difference sign between the two compounds [22]. We obtained the experimental node-level ground truths for attributions, and the edge attributions can be equally redistributed onto their endpoint node attributions.
In this paper we tested multiple feature attribution methods that enable the estimation of positive and negative atom contributions [33] including Class Activation Maps (CAM) [34], Gradient $\times$ Input (DeepLIFT)[26], Gradient-weighted Class Activation Mapping (Grad-CAM) [27], and Integrated Gradients (IG) [35]. For CAM, a global average pooling (GAP) layer takes node and edge activations and sums them to create a graph embedding layer before outputting to obtain attributions [34]. Gradient $\times$ Input computes attributions by taking the input graph’s element-wise product and the predicted output’s gradient for the input node and edge features. It can also include a reduction step across the feature dimension to produce node- or edge-level attributions [26]. Grad-CAM utilizes the gradient of predicted output for the intermediate activations representing a transformed version of the input to measure input importance and remove the necessity of a GAP layer in CAM [27]. IG integrates the element-wise product of an interpolated input with the gradient of the predicted output for the interpolated input, between the actual input and a counterfactual input [35].
2.4 Evaluation & Explanation metrics
We utilized the root mean squared error (RMSE) and Pearson’s correlation coefficient (PCC) metrics of predicted vs. true binding affinity IC50 to estimate model performance against three kinases respectively. Although many classification models [23] exist for drug property classification, such as predicting the presence of toxicity and molecules as kinase inhibitors, there are still limited prediction models that maximize the use of molecular structural information and mitigate the impact of training set imbalance to predict drug affinities like IC50 on a specific target protein. Moreover, classification labels, such as identifying potential molecules as kinase inhibitors or not, are often determined based on arbitrary thresholds of drug properties like IC50. Consequently, we selected the benchmark model trained with the uncommon-node-focused loss function proposed by Kenza Amara et al. [31]. We then compared it against our models, which use loss functions that integrate both common and uncommon node information under various regularization schemes (see Table 1). They also introduce global direction as a binary metric assessing whether average feature attribution across the uncommon nodes in a pair of compounds preserves the direction of the activity difference. We evaluated the performance of feature attribution methods by using global direction. Assuming $\psi:C\rightarrow\mathbb{R}^{N\times d}$ be a feature attribution function that assigns a score to each node feature in an input graph and the rest of function settings were the same as the loss functions including the readout functions $R(\cdot)$ and the masking function $F(\cdot)$, the computed metric for one pair $m_{k}=(c_{i},c_{j})$ is:
$g_{dir}(m_{k})=$ $\mathbbm{1}[\text{sign}(R_{i}(F_{i}(\psi(c_{i}))-R_{j}(F_{j}(\psi(c_{j})))=\text{sign}(y_{i}-y_{j})].$
Furthermore, we compared averaged global direction scores averaged across three kinases by the Wilcoxon signed-rank test [36] based on the loss function ($L_{MSE}+L_{N}$) with and without group lasso. The scores for each feature attribution method were averaged over compound pairs in the test sets (see Section 3.2 for details). We also considered the instability of atom-level accuracy, which measures whether the feature attribution assigned to a node has the same sign as the experimental activity difference of the compound pair (see Section 3.3 for details).
For model explanation, we assumed that the mentioned loss functions designed for GNN models would contribute to the graph-level prediction of the molecular property, and statistical regularization methods could prune and highlight vital molecular substructures at the atom-level molecular property prediction. We first performed explainability evaluation at the graph level by using the global direction scores, which capture both common and uncommon nodes for a compound pair. Thus, we can assess whether the direction of the activity difference is preserved. We then conducted explainability for individual protein targets by mapping feature attribution values to two compound structures in the context of their binding receptors. An increase in feature attribution performance reflects an improvement in model explainability.
3 Results
3.1 Model performance
We assessed model performance under different combinations of loss functions and regularization schemes via 5-fold cross validation. Table 1 compares model prediction performance by using averaged root mean square error (RMSE) and Pearson correlation coefficient (PCC) values and their weighted values based on the testing set. We evaluated RMSE and PCC on all three test datasets, first by assigning equal weight to each dataset and then by weighting them according to the count of active-cliff pairs in their respective test sets. We defined the MSE loss on the absolute predicted versus experimental binding affinities of molecule pairs, increased loss function complexity from only uncommon node information to both common and uncommon substructures, and added group lasso or sparse group lasso regularization to such loss functions. They have an increasingly positive influence on model performance, reflected by the decreasing trend of averaged RMSE values and the increasing trend of averaged PCC values. The model with the loss function considering both common and uncommon substructures and applying sparse group lasso performs the best compared with other loss function settings, which achieves the lowest averaged $\text{RMSE}=0.2143$ and lowest weighted averaged $\text{RMSE}=0.1984$, and the highest averaged $\text{PCC}=0.9615$ and lowest weighted averaged $\text{PCC}=0.9468$. Moreover, drug property prediction was improved by integrating common and uncommon node information and using regularization, reducing the averaged RMSE by 15.71% and the weighted averaged RMSE by 11.62% (with group lasso), and decreasing the averaged RMSE by 17.48% and the weighted averaged RMSE by 11.62% (with sparse group lasso) compared to the loss function only considering the uncommon nodes without penalties.
Table 1: Model performance of different loss functions (5-fold cross-validation)
[…]
$L_{\text{UCN}}$a 0.2597 ($\pm$0.0254) 0.9541 ($\pm$0.0096) 0.2245 ($\pm$0.0303) 0.9384 ($\pm$0.0164)
$L_{\text{N}}$b 0.2371 ($\pm$0.0221) 0.9564 ($\pm$0.0056) 0.2143 ($\pm$0.0278) 0.9386 ($\pm$0.0161)
$L_{\text{N+GL}}$c 0.2189 ($\pm$0.0219) 0.9609 ($\pm$0.0071) 0.1984 ($\pm$0.0205) 0.9468 ($\pm$0.0146)
$L_{\text{N+SGL}}$d 0.2143 ($\pm$0.0185) 0.9615 ($\pm$0.0078) 0.1984 ($\pm$0.0150) 0.9466 ($\pm$0.0140)
  • $\mathrm{\dagger\ddagger}$
    Averaged root mean square error (avg. RMSE) and Pearson correlation coefficient (avg. PCC) across three datasets equally, and weighted average using their number of active-cliff pairs in testing sets (w. RMSE & w. PCC)after 5-fold cross-validation.
  • a
    Loss function only for uncommon substructure (same setting as Kenza Amara et al. [31]).
  • b
    Loss function for both common and uncommon substructure.
  • c
    Loss function with group lasso.
  • d
    Loss function with sparse group lasso.
To demonstrate the consistency of model performance, we also compared the performance of models targeting three different kinases in the testing sets respectively under various loss function settings (Figure 3). The results indicate consistent trends of decreasing RMSE and increasing PCC as more node information is included, progressing from only uncommon nodes to both common and uncommon nodes. Additionally, increasing the complexity of the model by adding more regularization items, such as moving from no penalties to including group lasso and sparse group lasso, leads to better model performance. This demonstrates that models with regularization outperform those without.
Figure 3: Model performance for the molecules in testing sets targeting three kinases respectively, under different loss function settings via 5-fold cross-validation. Numbers for each bar plot indicate the averaged RMSE or PCC values for each loss function setting, and error bars represent standard deviation. It shows the consistent trends of RMSE decreasing and PCC increasing as more node information (from only uncommon nodes to both common and uncommon nodes) and regularization methods were added (from no penalty items for loss functions to adding group lasso and sparse group lasso). This suggests that the models using loss functions with regularization perform better on affinity prediction.
3.2 Model Evaluation
We evaluated and compared the averaged global direction scores for $L_{MSE}+L_{N}$ with and without group lasso by scatter plots with connecting lines for the four feature attribution approaches respectively, using compounds targeting three kinases in the testing sets. The used compound pairs are considered at the minimum 50% MCS threshold from 50% to 95% in 5% increments. Figure 4 suggests a relationship where the global direction scores predicted by the model with the loss function incorporating group lasso (y-axis) tend to exceed those of the model without the penalty item (x-axis) across the testing dataset under different minimum common substructure thresholds. It shows that there is an increase of 47.13% in the global direction score for the feature attribution CAM with a significant Wilcoxon test p-value of 0.0002 (Figure 4A), an increase of 14.83% in the global direction score for Grad-CAM with a significant Wilcoxon test p-value of 0.0059 (Figure 4B), an increase of 15.4% in the global direction score for Gradient $\times$ Input with a significant Wilcoxon test p-value of 0.002 (Figure 4C), and an increase of 8.49% in the global direction score for IG with a significant Wilcoxon test p-value of 0.0098 (Figure 4D). Supplement figures (Figure 6, 7, and 8) also include scatter plots with connecting lines for compounds targeting three kinases respectively, reflecting the increase of global direction scores for $L_{MSE}+L_{N}$ with group lasso compared to the ones without group lasso. The results suggest that introducing regularization like group lasso improves the performance of feature attribution methods.
Figure 4: Comparison of averaged global direction scores for $\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{N}}$ with and without group lasso by scatter plot with connecting lines. The plot was used for comparing graph-level global direction scores to show the distribution difference of averaged predicted global direction values with $\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{N}}$ (x-axis) and $\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{N}}$ with group lasso (y‑axis) loss functions under different minimum common substructure thresholds. Compound pairs are considered at the minimum 50% MCS threshold and from 50% to 100% in 5% increments. The text box reports the increased percentage of global direction scores with group lasso regularization for all the feature attribution including CAM (Figure. 4A) having 47.13% increase and a significant Wilcoxon test p-value of 0.0002; Grad-CAM (Figure. 4B) having 14.83% increase and p-value of 0.0059; Gradient $\times$ Input (Figure. 4C) having 15.4% increase and p-value of 0.002; IG (Figure. 4D) having 8.49% increase and p-value of 0.0098.
3.3 Exemplary explanations for molecules in test set
We compared the atom-level accuracy in atom coloring for ground truth feature attribution labels, the atom coloring using the feature attribution approach Grad-CAM with MSE loss and node loss without penalty, and the one using Grad-CAM with two loss functions with sparse group lasso. Figure 5 shows the structural explanations of one example ligand binding to the kinases 1O42, 2H8H, and 4MXO respectively from the testing sets. In Figure 5, the first column shows the ground truth feature attribution, while the latter two columns present the atom coloring captured and predicted by the feature attribution method Grad-CAM with loss function $L_{MSE}+L_{N}$ with or without penalties. It shows that the predicted coloring of nodes is more consistent for the model trained by using the loss function $L_{MSE}+L_{N}$ with penalty than the model with non-penalty settings. The results suggest that utilizing regularization, such as sparse group lasso, improves model explainability for graph-level drug property prediction and the identification of vital molecular substructures.
Figure 5: Comparison of the atom-level accuracy in node coloring for ligands binding to the three kinases. It shows the ground truth feature attribution labels of atom coloring and prediction under three situations: feature attribution approach Grad-CAM with MSE loss and node loss functions without penalty, and the one using Grad-CAM with the two loss items with sparse group lasso. When the sparse group lasso was applied for the loss functions, the predicted atom coloring was much more consistent with the ground-truth feature attribution labels of coloring.
4 Discussion
This paper provides a computational and statistical framework for compound-protein affinity prediction by utilizing the activity-cliff data, exploiting both common and uncommon node information to train GNN models, and determining vital substructures by regularization methods such as group lasso and sparse group lasso. It shows the improvement of GNN model performance by using GNN for capturing and integrating common and uncommon substructure information of activity-cliffs pairs and relying on loss functions with regularization to prune and highlight molecular subgraphs. It also demonstrates the improvement of global direction scores and atom-level accuracy in atom coloring predictions by applying regularization methods to enhance feature attribution, thereby providing stable model explainability for molecular property prediction. Moreover, this approach can potentially enhance the interpretability of models in drug discovery and virtual screening pipelines, especially when investigating crucial molecular substructures during lead optimization.
Due to limited time and resources for model training and parameter tuning with different loss function settings, we tested and verified the methods on the three Src kinases and compared the performance of feature attribution and model settings. Src has been identified as a promising target for Alzheimer’s disease (AD) therapeutics involved in many AD progression pathways [17]. In addition, oncogenic activation of Src family tyrosine kinases has been demonstrated to play an important role in solid cancers, promoting tumor growth and the formation of distant metastases [18]. Furthermore, due to its importance, there are abundant compound pairs (at least 50 pairs) in the training set for them. Utilizing compounds targeting these three kinases may mitigate the impact of data imbalance, where inactive compounds outnumber active ones. However, it may also introduce another form of data imbalance, where compounds sharing common scaffolds and motifs with many other compounds appear in numerous activity-cliff pairs, potentially affecting model training and prediction performance. Verifying the generality of the model’s predictive performance will require extensive computational resources and time for model training with more activity-cliff pairs targeting a broader range of proteins.
Even though this method implements regularization approaches such as group lasso and sparse group lasso to prune and select important common and uncommon node substructures, an average pooling layer was used to normalize the node information during model training. It may weaken the signals from vital nodes and their substructures and reduce the model’s sensitivity to identify important molecular substructures. This is reflected by the lighter and more even node coloring prediction compared to the ground-truth feature attribution in Figure 5. If we hope to compute the loss that correlates decoration embedding and activity cliffs at individual decoration sites, we can further define such loss functions based on individual node information.
Acknowledgment
The authors would like to thank the support from the Target Enablement to Accelerate Therapy Development for Alzheimer’s Disease (TREAT-AD) drug discovery center at Indiana University School of Medicine.
Supplementary
S1. Global direction scores for test-set molecules targeting 1O42
Figure 6: Comparison of global direction scores by using test-set molecules targeting 1O42 for $\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{N}}$ with and without group lasso by scatter plot with connecting lines. The text box reports the increased percentage of global direction scores with group lasso regularization for all the feature attribution including CAM (Figure. 6(a)) having 51.37% increase and p-value of 0.0371; Grad-CAM (Figure. 6(b)) having 28.07% increase and p-value of 0.002; Gradient $\times$ Input (Figure. 6(c)) having 17.03% increase and p-value of 0.002; IG (Figure. 6(d)) having 7.9% increase and p-value of 0.0547.
S2. Global direction scores for test-set molecules targeting 2H8H
Figure 7: Comparison of global direction scores by using test-set molecules targeting 2H8H for $\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{N}}$ with and without group lasso by scatter plot with connecting lines. The text box reports the increased percentage of global direction scores with group lasso regularization for the feature attribution including CAM (Figure. 7(a)) having 40.06% increase and and p-value of 0.0075; Grad-CAM (Figure. 8(b)) having 12.4% increase and p-value of 0.0075; Gradient $\times$ Input (Figure. 7(c)) having 61.68% increase and p-value of 0.0075; IG (Figure. 7(d)) having 8.56% increase and p-value of 0.0075.
S3. Global direction scores for test-set molecules targeting 4MXO
Figure 8: Comparison of global direction scores by using test-set molecules targeting 4MXO for $\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{N}}$ with and without group lasso by scatter plot with connecting lines. The text box reports the increased percentage of global direction scores with group lasso regularization for the feature attribution including CAM (Figure. 8(a)) having 69.62% increase and p-value of 0.002; Grad-CAM (Figure. 8(b)) having 18.27% increase and p-value of 0.0077; Gradient $\times$ Input (Figure. 8(c)) having 9.96% increase and p-value of 0.002; IG (Figure. 8(d)) having 10.74% increase and p-value of 0.00109.